OpenStudy (anonymous):

Can someone explain to me how you solve a problem like this one below? Will medal and follow.

OpenStudy (anonymous):

Determine which two values the following expression is between. \[2\sqrt{41}\] 12.7 and 12.8 13 and 13.1 12.9 and 13 12.8 and 12.9

OpenStudy (anonymous):

Use your calculator to approximate \( 2 \sqrt {41}\)

OpenStudy (anonymous):

it it 12.8 and 12.9

OpenStudy (anonymous):

how would I use my calculator for it? Im not sure which button to use.

OpenStudy (anonymous):

\[ 2. \sqrt{41}\approx 12.8062 \]

OpenStudy (anonymous):

2 times the sqrt of 41

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (anonymous):

does your calculator has a sqrt button

OpenStudy (anonymous):

Yeah it does, I get how to do it now

OpenStudy (anonymous):

Excellent